OpenStudy (anonymous):

Tom can fix a car in 6 hours. Ellen can do it in 4 hours. How long would it tak them working together?

OpenStudy (phi):

one way to think of these problems is rate * time = job (almost like rate * time = distance) tom's rate is 1 job/6 hours (1 car per 6 hours) ellen's rate is ¼ if you add them: (1/6 + ¼) t = 1 can you solve for t ?

OpenStudy (anonymous):

I would add 1/6 + 1/4? @phi

hero (hero):

Let T be how fast Tom can fix the car alone Let E be how fast Ellen can fix the car alone Let x be how fast Tom and Ellen can fix the car together Then \[\frac{T \times E}{T + E} = x\] Since in this case T = 6 E = 4 Then \[\frac{6 \times 4}{6 + 4} = x\]

OpenStudy (anonymous):

24/10?

hero (hero):

or 2.4 Hours in decimal form

OpenStudy (anonymous):

Ok the answer in the book is 2 hrs 24 minutes. How did they get that?

hero (hero):

Because they converted 4/10 of an hour to minutes

hero (hero):

.4 = 4/10

hero (hero):

You can set up a proportion such as this to find minutes: \[\frac{4}{10} = \frac{x}{60}\]

hero (hero):

If you have the fraction of hours, you can convert it to minutes by setting the hour fraction equal to the minutes fraction x/60

hero (hero):

Then solve for x

hero (hero):